Download
muon operations summary n.
Skip this Video
Loading SlideShow in 5 Seconds..
Muon Operations Summary PowerPoint Presentation
Download Presentation
Muon Operations Summary

Muon Operations Summary

152 Views Download Presentation
Download Presentation

Muon Operations Summary

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Muon Operations Summary A. Polini (on behalf of the Muon Group) Outline • Summary from Operations Session • Shutdown Activities • CSC, MDT, RPC, TGC • MDT EE Side C • Readiness for 2012 • First Running • Data taking • Plans • Manpower • Shift Organization RPC CSC MDT TGC Many thanks to all who provided material and information

  2. CSC CSC Operation New ROD firmware V08-00-00● Removal of status words● Time window ● Two-sample readout ● ROD register checks Fixed broken fiber couplerChannel status Michael Schernau Slide 1 0

  3. CSC Removal of status words Old release V070603  37% busy at 75kHz } expected data volume New Release V080000  ~5% busy at 75kHz } expected data volume  Removed status words for error-free events No impact on physics or debugging Slide 2 03/19/12

  4. CSC Two sample readout New option to read only 2 instead of 4 samples Reduces clusters from~10 words to ~6 words Data volume reduction } expected data volume  3.6 kB--> 2.0 kB  5.7 kB--> 3.5 kBReduces busy 4% --> 1.5% 8% --> 3.2% Slide 3 03/19/12

  5. CSC Four- and two-sample readout Four sample readout  Resolution 90 microns 80% unspoiled hits Timing-independent amplitudes Two sample readout Resolution 98 microns 2011 Z --> mu mu data  83% unspoiled hits - Due to slightly smaller amplitudes  No more time info Slide 4 03/19/12

  6. CSC Smaller readout time window Current time window is ~140 ns wide Optional new time window of 75 ns width  Reduces hits from previous and next bunch crossing Reduces hits from early beam halo  Expected benefit: 30% reduction in data volume Requires the installation of fiber extensions to equalize timing  The window can be moved in 25 ns time intervals - If not properly aligned, efficiency losses for early orlate hits can occurr Initial tests show that the execution time increasesbusy by 0.5 % which would reduce the gain fromdata volume  Slide 5 03/19/12

  7. CSC Deadtime versus L1 rate after busy ATLAS-wide test, but no occupancy in other detectorsSimulate L=7e33, two-samples, 75ns time window Slide 6 03/19/12

  8. CSC ROD register monitoring Internal registers of the ROD are read● During configuration ● Before start of run● At end of run Checked against expected valuesErrors are sent to MRS and log filesTransfers took too long ● Commands were serviced every 0.5 seconds 14 commands took 7 seconds ● One register had a link-lock related flag set ● This flag is always set and is now ignored Slide 7 03/19/12

  9. CSC Fixed broken fiber coupler Consistent link loss during HI run Unlocked fiber producesrandom data High occupancy in sectorA05 layer 1 and 2 Fixed by replacing the fiber connector of the control fiber at the patchpanel Slide 8 03/19/12

  10. CSC Detector status Unchanged from lastyear 3 layers without HV 32 groups of 12 channelswith bit errors 96.4% OK SideC SideA Slide 9 03/19/12

  11. CSC Summary Removal of status words  37%  5%, no impact on physicsTwo sample readout  5%  ~3%, resolution worse by 8 microns, loss oftiming information Even almost empty events give ~1% busy at 75kHz 2-4% busy will be unavoidable  Busy decreases when L decreases during run 96.4% of all channels are OK Ready for 2012 data taking Slide 10 03/19/12

  12. MDT MDT Pit Activities Bing Li Univ. of Science & Tech. of China Institute of Physics, Academia Sinica On Behalf of MDT Group Muon Week MeetingMarch 19th, 2012

  13. MDT MDT Installation and Maintenance • EE chamber installation, services and commissioning  Installation by CERN team;  Chamber survey and Alignment service Gas Services  Cable Services, HV/LV and readout mapping Commissioning • Gas Maintenance  Zeolith gas filter replacement (Joerg + MPI Technicians) Gas leak test and measurement (Joerg) • EO Gas Leak Repair (Anatoli + UM + USTC personals) • MDT Chamber Maintenance  Repair damaged chambers during shutdown activities;  Replacing front-end electronics to fix issues on readout;  Resolving B-field, temperature and voltage sensor problems; • Power Suppliers (George, Stephanie …, detail see Preema Talk) 2

  14. MDT EE Chamber Installation, Services and Commissioning 3

  15. MDT Manpower for Services => Installation CERN team, Giancarlo Spigo etc. => Gas Joerg Dubbert, Jianbei Liu => Cable Service installation at pit Claudio Ferretti, Zhen Yan, Anatoli Kozhin, Lulu Liu, Bing Li, Tiesheng Dai => Alignment Christoph Amelung, Laurel Coffey, Benjamin Wiener, Andrei Dushkin, Jim Bensinger, Kevan Hashemi and Survey team from CERN => Technicians Dmitri Kharchenko, Konstantin Levterov, Nikolay Kusnetso 4

  16. MDT Manpower for Commissioning => Detector Operation and Commissioning from P1 Control Room Jacob Groth-Jensen, Silvia Franchino, Rosemarie Aben, Alan W Wilson, Yusheng Wu, Laura La Rotonda, Lulu Liu, Zhen YanTiesheng Dai, Georgios Iakovidis, Konstantinos Ntekas, Konstantinos Karakostas, Stefanos Leontsinis => Supports ELMB and MROD: Robert Hart, Henk Boterenbrood, GerjanBobbink LV/HV/DCS Control: Stephanie Zimmermann, Georgios Iakovidis, Konstantinos Ntekas, Konstantinos Karakostas, Stefanos Leontsinis TDAQ: Enrico Pasqualucci, Preema Rennee Pais GNAM: Gaudio Gabriella GNAMON: Philipp Fleischmann Others: Devin Harpe (MDT Config. DB), Hugh Skottowe 5

  17. MDT Few unexpected issues for EE Installation at Beginning - C07 alignment bar too long - EEL mounted in moving rack wrongly, ML1 was facing out instead of ML2, resolved - EELC11 could not installed since crane was blocked by endcap toriod magnet, resolved First Installed EE chamber: EEL1C09on Jan. 6th, 2012 6

  18. MDT Commissioning Goals • Functional alignment system and consistent results with survey, and adjust chamber to final position (Christoph) • No gas leak (Joerg and Jianbei) • CANBUS and DCS functional • Front-end electronics hold LV and make sure JTAG initialization works. • Continues monitor B-field, Temperature and Voltage Sensors to ensure reasonable readout • Chamber hold HV 3080V with low dark current • Right mapping for LV, HV and readout • No unexpected noise • No unexpected dead tube/channel • Reasonable ADC (peaked around 100ns) and flat TDC distributions) Noise rate of all tube/channel, all below 5KHz at threshold of -50mV Noise rate of all tube/channel, allbelow 5KHz atthreshold of -50mV 8

  19. MDT Alignment • Route and Connect Alignment cables • Survey sector • Readout alignment sensor data followed by analysis • Adjust chamber(s) if necessary (typically by few cm) • Readout alignment to verify • After fixing issues on 3 alignment sensors, all of them works now; • Need to cut access platforms for EES to link different EE sectors since they block optical path where most of them already done; • Still need to install optical sensors for sector C05 and replacing cable during next technical stop . 9

  20. MDT EE alignment system works, still working on analysis to get a full picture ofEE alignment, details see Wed. Christoph talk on Endcap Muon Alignment. Just show one example: Adjusted EEL-C Position 10

  21. MDT EE Gas Services • Put gas connector on gas pipe and install extension pipe if necessary before EE chamber/sector installation; • Check chamber pressure and depressurize it after installation; • Wait till final chamber adjustment done; • Connect to gas system • Check the gas leak (~1 day) • Gas flush at 2 bar for about 1 day • Pressurize to 3 bar No major leak, all of them less than few mbar (limited bymeasurement precision), except of EES1C08 ML2 ~20mbar per day 11

  22. MDT EE Cable Service Sequence • Install HV splitter boxes for EEL chambers • Install and connect the safety grounding cables from chamber to general copper grounding line • Route DCS cables to EE chambers and connect to corresponding MDT-DCS box (MDM) • Route LV and HV cables to EE chambers and connect to corresponding CSM and chamber multilayer • Install and route readout fibers from fanout box to EE chambers and connect to corresponding CSM and fiber fanout channel, where the fiber fanout box on sector 9 need to be installed Turn on both ELMB and CSM LV to check DCS fully functional and no problem in JTAG initialzation 13

  23. MDT B-field Sensor Measurements B-field : 0.5 to 1.5G Temperature: around 21.5~23 ℃  All fine 14

  24. MDT Chamber Temperature Measurements Example: EES1C06 All 8 temperaturesensors have goodmeasurements.  Similar for otherchambers, all works 15

  25. MDT Front-end Electronics T & V Sensors Example: EEL1C09 All V measurementsas expected, and alltemperature value are reasonable All chambers are OK 16

  26. MDT LV and HV Mapping LV Mapping: Turn on sector by sector following the cable service ensure right LV mapping both in LV cable connection and DCS control HV Mapping with 2 steps: 1) Measure HV line by line (i.e., ML by ML) with 30V at chamber end to ensure right HV connection; 2) Collect data by turning on only 1 ML and check ADC distributions which ensure right mapping in DAQ and monitor Charge Distribution: Charge Distribution: Cosmic Hits on ML2 No Cosmic Hits on ML1 (HV only on ML2) 17

  27. MDT Fixed EE Hardware Problems Even though EE chambers have been fully tested, See Lulu’s talk https://indico.cern.ch/conferenceDisplay.py?confId=162647 still one has to: • Replace C09 fiber fanout due to one bad TTCchannel; • Replace one broken fiber from fiber fanout box toMDT chamber • Replace CSM card for EES2C12 (twice) and EES1C10 • Replace mezz card 4 for EES2C12 to fix unexpecteddead tubes due to loose connector • Disconnect EES1C10 ML2-2-22 from HV 20

  28. MDT EEL A03 Installation and Service Only giving half day for service works • EEL A03 installation done around noon of Feb. 21,2012; • A03 survey and alignment services done afternoon • Obtain initial alignment results early evening • Gas services finished at night • Cable services finished at night • DCS, JTAG and readout fully functional at same day • Final alignment, gas and HV were fullycommissioned on Feb. 22 before 14:00  We did it 21

  29. MDT Check All Install EE side A fibers GOL TTC GOL TTC -25.38 -24.74 EEL1A01 -25.01 -24.85 EEL2A09 EEL2A01 -25.09 -24.89 EES1A10 -25.38 -24.89 EES1A02 -25.34 -25.38 EES2A10 -25.69 -24.85 EES2A02 -25.64 -24.97 EES1A12 -25.25 -25.01 EEL1A03 -27.5 -24.93 EES1A12 -25.51 -25.42 EEL2A03 -25.46 -24.89 EES1A14 -26.02 -24.97 EEL1A07 -25.51 -25.09 EES2A14 -25.69 -25.21 EEL2A07 -25.74 -25.05 EEL1A15 -25.6 -24.85 EES1A08 -25.51 -24.93 EEL2A15 -25.33 -24.78 EES2A08 -25.55 -25.05 EES1A16 -25.55 -24.38 EEL1A09 -25.64 -24.66 EES2A16 -25.38 -24.35 All fine 22

  30. MDT MDT Maintenance • ~30 problems were investigated from gas leak, HV trip to various sensor issues; • Fixed 2 gas leaks; • Fixed 3 HV trips; • Replace 1 ELMB; • Replace 5 CSMs; • Replaced ~10 mezzanine cards; • Replaced 1 B-field sensor; • Two broken mezzanine cables (someone stepped on them!), one fixed and one still need to be done properly; • Touch/push/reconnect mezzanine cables. 1. Most mother boards are EU-MB-04-21-xxxx except two cases with EU-MB-04-20-xxxx 2. At least 4 cases, mezzanine card got JTAG problem where downstream mezzanine will be killed if troubled mezzanine is included. 27

  31. MDT MDT Conclusion • EE chambers of full side C and A03 were installed and commissioned in short period; • New installed EE chambers are gas tight, low noise, fully functional and integrated into Muon/ATLAS TDAQ, all B-fieldand Temperature sensors work well; • A major effort is done to fix EO gas leak; • MDT maintenance are performed and many problems are fixed. 28

  32. MDT • EE integration • DCS/TDAQ • Services Gas • Manpower A. Polini

  33. MDT A. Polini

  34. MDT A. Polini

  35. MDT Summary (MDT Operations) • New EE chambers included in DCS, TDAQ configurations • Many DCS updates completed during shutdown • TDAQ migration to v-04-00-01 went smoothly, only a few minor issues to be fixed with MDT DAQ • Manpower for DCS and TDAQ (on-call experts) is sufficient for now, but replacements may be needed by the end of the year - training is needed for new people A. Polini

  36. TGC

  37. TGC

  38. TGC

  39. TGC

  40. TGC

  41. TGC

  42. TGC

  43. RPC

  44. RPC

  45. RPC A. Polini, S. Zimmermann

  46. RPC

  47. RPC

  48. RPC

  49. RPC

  50. RPC